# LRFusionPR: A Polar BEV-Based LiDAR-Radar Fusion Network for Place Recognition

### v2

@@ III-D Network Training
$\phi_{\text{F}}=\langle\frac{\mathcal{D}_{\text{F,pos}}-\mathcal{D}_{\text{F,q}}}{||\mathcal{D}_{\text{F,pos}}-\mathcal{D}_{\text{F,q}}||_{2}},\frac{\mathcal{D}_{\text{F,neg}}-\mathcal{D}_{\text{F,q}}}{||\mathcal{D}_{\text{F,neg}}-\mathcal{D}_{\text{F,q}}||_{2}}\rangle$ (11)
$\phi_{\text{A}}=\langle\frac{\mathcal{D}_{\text{A,pos}}-\mathcal{D}_{\text{A,q}}}{||\mathcal{D}_{\text{A,pos}}-\mathcal{D}_{\text{A,q}}||_{2}},\frac{\mathcal{D}_{\text{A,neg}}-\mathcal{D}_{\text{A,q}}}{||\mathcal{D}_{\text{A,neg}}-\mathcal{D}_{\text{A,q}}||_{2}}\rangle$ (12)
where $\langle$,$\rangle$ , ⟩ represents the inner product operation.
@@ IV-A Datasets and Experimental Setup
To comprehensively validate our proposed method, we conduct experiments on three autonomous driving datasets: nuScenes [31], MulRan [12], and Oxford Radar [29].
nuScenes. The dataset features diverse weather, illumination changes, and multiple revisits, with 32-beam LiDAR and single-chip radar data. Following AutoPlace [4], we train and evaluate the model on the most comprehensive split, BostonSeaport (BS). In addition, we validate the generalization ability on splits collected from unseen regions, namely SingaporeOneNorth (SON) and SingaporeQueenstown (SQ). For the SON and SQ splits, we partition the data similarly to [4], taking the data from the first 30 days and 15 days as the database set and the rest as the test query set.
MulRan. The dataset emphasizes long revisit intervals and significant viewpoint transformations, with 64-beam LiDAR and scanning radar data. We perform place recognition evaluation on the Sejong, Riverside and DCC splits. We organize the Sejong split according to [5], and the DCC split following [32]. For the Riverside split, we use session 01 as the database set and session 02 as the test query set. Due to the lower sampling frequency of the scanning radar, we discard LiDAR data that cannot be time-synchronized with the radar to ensure multimodal data consistency.
Oxford Radar. Collected along a specific route in Oxford, the dataset includes 32-beam LiDAR and scanning radar data with temporal variations. We select the session ”2019-01-11-13-24-51” as the database set. The session ”2019-01-16-14-15-33”, collected in rainy weather, is used as the test query set. We choose these sessions to evaluate the robustness of place recognition under adverse weather conditions. More details of our training and evaluation splits are shown in Tab. I.
@@ IV-B Implementation Details
We set the size of the LiDAR polar BEV to $h_{\text{P}}=200$ and $w_{\text{P}}=900$, and the radar polar BEV to $h_{\text{R}}=50$ and $w_{\text{R}}=225$. The maximum perception range $m$ is set to 80 m. For the NuScenes radar data, we stack seven nearest frames before projecting them into polar BEVs following [4]. We apply height-based ground filtering to process LiDAR data. For the polar cross-attention module, we set the feature embedding dimension $d_{\text{model}}=128$, the feed-forward dimension $d_{\text{ffn}}=1024$, and the number of heads $n_{\text{heads}}=8$. For the NetVLAD module, we set the number of clusters $d_{\text{clusters}}=64$, and the dimension of the descriptor $N_{\text{D}}=256$. The final descriptor then has a size of $1\times 512$. We set the margin of the Lazy Triplet Loss to $\alpha=0.5$, the margin of the structure-aware loss to $\beta=0.0002$, the weighting parameter $\lambda=5$, and the number of negative samples to $k=10$. We optimize the model using the Adam optimizer with an initial learning rate of $5\times 10^{-5}$. The learning rate is halved every 5 epochs using the StepLR strategy. We apply random translation and rotation for data augmentation. When training on MulRan [12] dataset, we additionally use the distillation weight decay strategy from [28]. All of our experiments are conducted on a system with an Intel i7-14700KF CPU and an NVIDIA GeForce RTX 4060Ti GPU.
@@
Sessions Train Test
Database Query Database Query
nuScenes-BS 6312 7075 6312 3696
nuScenes-SON 0 0 4007 1184
nuScenes-SQ 0 0 418 851
MulRan-Sejong 14327 14327 1505 1380
MulRan-Riverside 0 0 1558 1724
MulRan-DCC 0 0 204 129
Oxford Radar 0 0 5657 6184
@@
C: Camera, L: LiDAR, R: Single-Chip Radar. The best and secondary results is highlighted in bold black and blue respectively.
@@
C: Camera, L: LiDAR, R: Scanning Radar. The best and secondary results is highlighted in bold black and blue respectively.
@@ IV-C Evaluation for Place Recognition
nuScenes results. Following [4, 9], we report Recall@1 (AR@1), Recall@5 (AR@5), Recall@10 (AR@10), and max $F_{1}$ score as metrics for measuring recognition accuracy. We compare LRFusionPR to state-of-the-art (SOTA) baselines, including unimodal (EgoNN [5], BEVPlace++ [7], AutoPlace [4]) and multimodal ones (MinkLoc++ [8], LCPR [9], CRPlace [24]). We train and evaluate baseline methods using their official open-source implementations. To ensure a fair comparison, we extend the maximum perception range of BEVPlace++ [7] from 40 m to 80 m to match other baselines. For CRPlace [24], we report the results from the original paper due to the lack of open-source code.
@@
As shown in Table II, our method demonstrates SOTA performance across various metrics on the AutoPlace [4] benchmark, as well as the unseen SON and SQ splits. By leveraging the complementary strengths of LiDAR and radar, our approach outperforms the LiDAR-only, radar-only and LiDAR-camera fusion baselines. In addition, we observe that LiDAR-camera fusion shows little performance improvement over LPR baselines. In the Autoplace benchmark, the database set contains only clear weather data, while the query set is collected in rain. We hypothesize that this may be due to model overfitting in training images.
@@
MulRan results. Since MulRan [12] does not include camera data, we use KidnappedRadar [3], EgoNN [5], and BEVPlace++ [7] as the baselines. The place recognition evaluation metrics follow the experiments on the nuScenes dataset. Notably, the higher sampling frequency of LiDAR enables LPR baselines to have more available training data than indicated in Tab. I. For fair comparison, we train BEVPlace and EgoNN following the data partition in [5], with 35.9k available LiDAR scans in the training set. In addition, we provide the evaluation results of the open-source EgoNN weights, denoted as $\text{EgoNN}^{{\dagger}}$, which is derived from joint training over multiple datasets.
@@
As shown in Tab. III, our method outperforms all baselines on Sejong split, and demonstrates superior generalization ability on Riverside and DCC splits. By leveraging LiDAR-radar fusion, our method effectively generates discriminative and robust multimodal place descriptors. This strikes a balance between accuracy and generalization.
TABLE IV: Comparison of cross-dataset generalization test
Methods AR@1 AR@5 AR@10 AR@20 max $F_{1}$
KidnappedRadar [3] 84.39 90.63 92.22 93.97 0.8450
EgoNN [5] 87.98 94.87 96.71 97.98 0.9367
BEVPlace++ [7] 94.36 97.51 98.57 99.05 0.9727
LRFusionPR (ours) 98.44 99.29 99.42 99.59 0.9931
@@
IV-D Study on Cross-Dataset Generalization
The MulRan [12] and Oxford Radar [29] datasets both include LiDAR and scanning radar data, but differ in sensor configurations and collection locations. These differences allow us to test cross-dataset generalization capabilities. We transfer pretrained weights of the baselines and our LRFusionPR from MulRan to the Oxford Radar dataset. As shown in Table IV, although the test query set from session “2019-01-16-14-15-33” is collected under heavy rain, our method still demonstrate superior recognition performance. By fusing LiDAR and radar modalities, LRFusionPR generates discriminative descriptors that achieve robust cross-dataset place recognition.
TABLE V: Place recognition accuracy under fog
Split Methods AR@1 AR@5 AR@10 AR@20
SQ AutoPlace [5] 64.12 66.12 68.00 72.24
MinkLoc++ [8] 23.15 46.89 58.28 69.92
LCPR [9] 13.98 27.03 37.49 49.24
EgoNN [5] 35.84 50.18 57.34 64.86
BEVPlace++ [7] 67.57 76.73 81.08 84.61
LRFusionPR (ours) 91.07 96.47 97.53 98.47
Sejong KidnappedRadar [5] 70.14 81.88 87.32 89.86
EgoNN [5] 40.07 55.14 61.67 66.74
BEVPlace++ [7] 56.16 70.94 76.74 81.96
LRFusionPR (ours) 93.48 98.41 99.42 99.86
TABLE VI: Ablation study on the effect of proposed component
@@ IV-D Study on Cross-Dataset Generalization
IV-E Robustness under Adverse Weather
In this section, we perform experiments on place recognition under adverse weather conditions. The robustness of LRFusionPR in rainy weather has been demonstrated in the nuScenes (Tab. II) and Oxford Radar (Tab. IV) experiments. We simulate fog scenarios on the SQ and Sejong splits to further validate the radar’s contributions. We follow the experimental design of MVDNet [25], using the fog model in [33] to simulate fog conditions. By adjusting the fog density parameter, we simulate fog with visibility ranging from 50 to 80 meters. In the experiments, we process all test query samples while keeping the database unchanged, simulating scenarios where vehicles encounter sudden weather changes. Fig. 4 illustrates the variation of Recall@1 with visibility, and  Tab. V shows the recognition accuracy at a visibility of approximately 50 m. Notably, the LiDAR-only and LiDAR-camera fusion baselines show significant performance degradation in foggy weather, whereas our LiDAR-radar fusion strategy demonstrates superior robustness. This shows that our method effectively exploits the weather resilience of both single-chip and scanning radars. Fig. 5 visualizes several simulated fog scenarios where BEVPlace++ [7] fails, while our method achieves correct recognition.
@@ IV-E Robustness under Adverse Weather
Figure 5: Visualization of place recognition in foggy weather. The green borders indicate correct retrieval results, while the red borders represent incorrect retrievals. We use red points to highlight the noise introduced by fog.
IV-F Ablation Study
In this section, we perform ablation experiments to investigate the effectiveness of our feature fusion mechanism. We analyze the impact of four key components on recognition performance: radar input (RA), polar cross-attention (PC), distillation branch (DB), and structure-aware loss (SA). For a comprehensive evaluation, we construct three types of datasets, each containing multiple splits, for the assessment of ablation study. We construct the single-chip radar dataset from nuScenes splits, and formulate the scanning radar dataset using MulRan and Oxford Radar splits. Furthermore, we integrate BS, Oxford Radar (rain), and SQ, Sejong (fog with visibility of 50 m) to create the adverse weather dataset. We compute the average AR@1, AR@5, AR@10 and max $F_{1}$ score over all splits in each dataset.
Our LiDAR-only baseline retains only the network architecture associated with the LiDAR input. As shown in Tab. VI (row 1), adverse weather severely degrades the recognition accuracy of it. Direct introduction of radar input with a simple concatenation-convolution fusion strategy fails to mitigate the sparsity and noise of the radar data, resulting in a loss of accuracy (row 2). Implementing polar cross-attention for multimodal fusion achieves an improvement, increasing robustness without sacrificing accuracy (row 3). Further integration of the radar-exclusive distillation branch significantly improves weather resilience while improving accuracy (row 4). Finally, we observe that optimization via structure-aware loss for self-distillation yields additional robustness gains. This suggests that cross-modality distillation enhances the discriminability of radar-exclusive descriptors, allowing them to play a more significant role in LiDAR degradation scenarios.
IV-G Study on Viewpoint Rotation Invariance

### v3

@@ III-D Network Training
[…]
where $\langle,\rangle$ represents the inner product operation.
@@ IV-A Datasets and Experimental Setup
To perform a comprehensive validation, we conduct experiments on four autonomous driving datasets: nuScenes [24], MulRan [25], Oxford Radar [26], and Boreas [27].
nuScenes. The dataset features rainy weather and multiple revisits, with 32-beam LiDAR and single-chip radar data. Following AutoPlace [4], we train and evaluate the model on the BostonSeaport (BS) split. In addition, we validate the generalization ability on unseen splits, namely SingaporeOneNorth (SON) and SingaporeQueenstown (SQ).
MulRan. The dataset emphasizes long revisit intervals and significant viewpoint transformations, with 64-beam LiDAR and scanning radar data. We perform evaluation on the Sejong, Riverside and DCC splits. We organize the Sejong split according to [5], and the DCC split following [28]. For the Riverside split, we use session 01 as database and session 02 as query. Due to the scanning radar’s lower sampling frequency, we discard LiDAR data not time-synchronized with the radar (time difference over 0.05 s) to ensure multimodal data consistency.
Oxford Radar. Collected along a specific route in Oxford, the dataset includes 32-beam LiDAR and scanning radar data with temporal variations. We select the clear session “2019-01-11-13-24-51” as the database set, and the rainy session “2019-01-16-14-15-33” as the test query set. We choose these sessions (OxfordRadar-Rain) to evaluate our method’s cross-dataset generalization capability in rainy conditions.
Boreas. The dataset provides real-world adverse weather data, including rain and snow, collected with a 128-beam LiDAR and scanning radar. We selected the clear-weather session “2020-12-18-13-44” as our database. The “2021-01-26-11-22” (Bor-Snow) and “2021-04-29-15-55” (Bor-Rain) sessions, collected during snowy and rainy weather respectively, served as query sessions. We utilize this dataset to demonstrate our method’s cross-dataset generalization capability under real-world adverse weather. More details of our training and evaluation splits are shown in Tab. I.
@@ IV-B Implementation Details
We configure the size of the LiDAR polar BEV as $h_{\text{P}}=200$ and $w_{\text{P}}=900$, and the radar polar BEV as $h_{\text{R}}=50$ and $w_{\text{R}}=225$, with 80 m perception range $m$. For the NuScenes radar data, we stack seven nearest frames before projecting them into polar BEVs following [4]. We apply height-based ground filtering to process LiDAR data. For the polar cross-attention module, we set the feature embedding dimension $d_{\text{model}}=128$, the feed-forward dimension $d_{\text{ffn}}=1024$, and the number of heads $n_{\text{heads}}=8$. For the NetVLAD module, we set the number of clusters $d_{\text{clusters}}=64$, and the dimension of the descriptor $N_{\text{D}}=256$. The final descriptor then has a size of $1\times 512$. We set the margin of the Lazy Triplet Loss to $\alpha=0.5$, the margin of the structure-aware loss to $\beta=0.0002$, the weighting parameter $\lambda=5$, and the number of negative samples to $k=10$. We optimize the model using the Adam optimizer with an initial learning rate of $5\times 10^{-5}$. The learning rate is halved every 5 epochs. We apply random translation and rotation for data augmentation. Given the more discriminative characteristics of scanning radar, we further adopt the distillation weight decay strategy from [22] when training on the MulRan dataset [25]. All of our experiments are conducted on a system with an Intel i7-14700KF CPU and an NVIDIA GeForce RTX 4060Ti GPU.
@@
Session Train Db/Query Test Db/Query Step
nuScenes-BS [24] 6312/7075 6312/3696 0.2 m
nuScenes-SON [24] 0/0 4007/1184 0.2 m
nuScenes-SQ [24] 0/0 418/851 0.2 m
MulRan-Sejong [25] 14327 1505/1380 0.2 m
MulRan-Riverside [25] 0/0 1558/1724 1.0 m
MulRan-DCC [25] 0/0 204/129 10.0 m
OxfordRadar-Rain [26] 0/0 6361/5657 0.2 m
Bor-Rain [27] 0/0 3386/3350 0.2 m
Bor-Snow [27] 0/0 3386/4327 0.2 m
@@
C: Camera, L: LiDAR, R: Single-Chip Radar. The best and secondary results are highlighted in bold black and underline respectively.
@@
C: Camera, L: LiDAR, R: Scanning Radar. The best and secondary results are highlighted in bold black and underline respectively.
TABLE IV: Place recognition performance under adverse weather conditions
Methods Modality Rain Simulated Fog Snow
OxfordRadar-Rain Bor-Rain SQ-Fog Sejong-Fog Bor-Snow
AR@1 AR@5 AR@10 AR@1 AR@5 AR@10 AR@1 AR@5 AR@10 AR@1 AR@5 AR@10 AR@1 AR@5 AR@10
AutoPlace [4] CR - - - - - - 64.12 66.12 68.00 - - - - - -
KidnappedRadar [3] SR 84.39 90.63 92.22 55.37 72.00 78.48 - - - 70.14 81.88 87.32 59.42 71.11 74.81
RadVLAD [16] SR 86.69 92.87 94.83 96.20 97.89 98.47 - - - 94.06 97.75 98.77 95.71 98.41 99.35
FFT-RadVLAD [16] SR 93.94 97.99 98.31 56.25 71.29 78.73 - - - 88.41 93.77 95.36 94.59 98.82 99.53
MinkLoc3Dv2 [12] L 92.68 97.35 98.32 94.36 98.24 99.28 75.91 91.42 94.24 65.72 81.67 87.17 19.51 32.77 38.66
EgoNN [5] L 87.98 94.87 96.71 98.81 99.67 99.82 35.84 50.18 57.34 40.07 55.14 61.67 71.13 84.28 87.89
BEVPlace++ [7] L 94.36 97.51 98.57 99.46 99.91 99.94 67.57 76.73 81.08 56.16 70.94 76.74 88.42 94.34 96.07
MinkLoc++ [8] L+C 93.07 97.19 97.88 85.67 96.45 97.97 23.15 46.89 58.28 - - - 3.51 8.32 14.01
LCPR [9] L+C - - - - - - 13.98 27.03 37.49 - - - - - -
LRFusionPR (ours) L+R 98.44 99.29 99.42 100.00 100.00 100.00 91.07 96.47 97.53 93.48 98.41 99.42 99.38 99.72 99.88
C: Camera, L: LiDAR, CR: Single-Chip Radar, SR: Scanning Radar, R: Compatible with both Single-Chip and Scanning Radar.
Missing entries in the table indicate that the session could not meet the input data requirements for the corresponding method.
@@ IV-C Evaluation for Place Recognition
nuScenes results. Following [4, 9], we report AR@1, AR@5, AR@10, and max $F_{1}$ score as metrics. We compare LRFusionPR to SOTA baselines, including unimodal (MinkLoc3Dv2 [12], EgoNN [5], BEVPlace++ [7], AutoPlace [4], RadVLAD [16], FFT-RadVLAD [16]) and multimodal ones (MinkLoc++ [8], LCPR [9], CRPlace [17]). We train and evaluate baselines using their official implementations. To ensure a fair comparison, we extend the maximum perception range of BEVPlace++ [7] from 40 m to 80 m to match other baselines. For CRPlace [17], we report the results from their paper due to the lack of open-source code.
@@
As shown in Table II, our method demonstrates SOTA performance across various metrics and sessions. By leveraging the complementary strengths of LiDAR and radar, our approach outperforms the baselines. In addition, we observe that LiDAR-camera fusion shows little improvement over LPR baselines. In the AutoPlace benchmark, where the database set contains only clear weather data and the query set is collected in rain, we hypothesize this is due to overfitting on training images.
@@
MulRan results. Since MulRan [25] does not include camera data, we use KidnappedRadar [3], RadVLAD [16], FFT-RadVLAD [16], MinkLoc3Dv2 [12], EgoNN [5], and BEVPlace++ [7] as the baselines. The place recognition evaluation metrics follow the experiments on the nuScenes dataset. For a fair comparison, we train BEVPlace++, MinkLoc3Dv2 and EgoNN using [5]’s data partition (35.9k LiDAR training scans). We also evaluate the open-source EgoNN weights, denoted as $\text{EgoNN}^{{\dagger}}$, which is derived from joint training over multiple datasets.
@@
As shown in Tab. III, our method outperforms all baselines on Sejong split, and demonstrates superior generalization ability on Riverside and DCC splits. This further demonstrates that our method effectively generates discriminative and robust multimodal place descriptors.
IV-D Robustness under Adverse Weather
In this section, we perform experiments on place recognition under adverse weather. To comprehensively evaluate the robustness of our proposed method in adverse weather conditions, we conduct experiments under three distinct weather types: rain, fog, and snow. Specifically, we select OxfordRadar-Rain and Bor-Rain as two rain-affected scenarios, and Bor-Snow for snowy conditions (see Sec. IV-A for details). Due to the lack of open-source fog datasets suitable for place recognition, we simulate fog conditions using the fog model from [29] following MVDNet [19]. We perform fog simulations on the Sejong and SQ sessions, which serve as our two fog test scenarios.
Tab. IV shows the recognition accuracy of our proposed method and the baselines across five scenarios representing different weather conditions. Additionally, Fig. 4 illustrates the variation of Recall@1 with visibility under simulated fog conditions. We find that LiDAR baselines suffer significant degradation in foggy and snowy conditions, where noise is prevalent and visibility is reduced. Although radar baselines demonstrate comparatively lower accuracy, they surpass LiDAR baselines in adverse weather. Notably, our method consistently outperforms both LiDAR-only, radar-only, and multimodal baselines across all adverse weather conditions, demonstrating the effectiveness of LiDAR-radar fusion for robust place recognition (see Fig. 5 for visualization). Furthermore, since our training set did not include Boreas [27] or Oxford Radar [26] data, these results demonstrate the cross-dataset generalization capability of our LRFusionPR.
@@ IV-D Robustness under Adverse Weather
TABLE V: Ablation study on the effect of proposed component
@@
TABLE VI: Ablation study on input modalities
Method SQ Sejong Riverside OxfordRadar-Rain SQ-Fog Bor-Snow
AR@1 AR@10 AR@1 AR@10 AR@1 AR@10 AR@1 AR@10 AR@1 AR@10 AR@1 AR@10
LRFusionPR-L 95.77 99.65 97.46 99.64 87.99 95.30 96.52 99.13 64.98 82.61 88.63 95.68
LRFusionPR-R 78.38 91.54 93.26 97.83 94.78 99.30 95.26 98.30 78.38 91.54 98.54 99.45
LRFusionPR 97.65 99.76 98.91 100.00 95.19 99.01 98.44 99.42 91.07 97.53 99.38 99.88
IV-E Ablation Study
In this section, we conduct ablation studies on the network architecture, input modalities, and the impact of multimodal training. We aim to demonstrate the rationale of the network design, and the contributions of each individual modalities for accurate and robust place recognition.
Network architecture. We analyze the impact of four key components: radar input (RA), polar cross-attention (PC), distillation branch (DB), and structure-aware loss (SA). For a comprehensive ablation, we construct three multi-split datasets: single-chip radar (from nuScenes), scanning radar (from MulRan and Oxford Radar), and adverse weather (integrating BS, OxfordRadar-Rain, SQ-Fog, and Sejong-Fog). We then compute the average metrics across all splits within each dataset. All different network architectures were trained from scratch. Our LiDAR-only baseline retains only the architecture associated with LiDAR input. As shown in Tab. V (row 1), adverse weather severely degrades its recognition accuracy. Direct introduction of radar input via a simple concatenation-convolution strategy failed to exploit discriminative radar features, leading to a loss of accuracy (row 2). Implementing polar cross-attention for multimodal fusion increases robustness without sacrificing accuracy (row 3). Further integration of the distillation branch significantly improves weather resilience (row 4). Finally, structure-aware distillation boosts the discriminability of radar-exclusive descriptors, yielding additional robustness gains. This validates the rationale of our network architecture design.
@@ IV-E Ablation Study
Figure 5: Visualization of place recognition in foggy weather. Each row in the figure corresponds to a different scene. The green borders indicate correct retrieval results, while the red borders represent incorrect retrievals. We use red points to highlight the noise introduced by fog.
Input modalities. To elucidate each modality’s role in our architecture, we conduct an ablation study on input modalities. We compare three input configurations: LiDAR-only (LRFusionPR-L), radar-only (LRFusionPR-R), and the full LRFusionPR. As shown in Tab. VI, while unimodal models can achieve comparable performance to the baselines, the superiority of LRFusionPR primarily stems from synergetic multimodal fusion. Even in clear weather with discriminative LiDAR data (SQ, Sejong), radar fusion yields further improvements. Meanwhile, in scenarios requiring long-range perception (Riverside), our approach effectively leverages the advantages of radar. In adverse weather (OxfordRadar-Rain, SQ-Fog, Bor-Snow), the LiDAR-only variant clearly degrades, especially in fog and snow. However, LRFusionPR fully exploits radar’s robustness, achieving robust place recognition through modal complementarity.
Impact of multimodal training. Here we validate the enhancement of radar modality through multimodal training. We compare the distillation branch trained with multimodal data (LRFusionPR-R-M) with that trained solely on radar data (LRFusionPR-R-S). As shown in Tab. VII, LRFusionPR-R-M outperforms its single-modality trained counterpart LRFusionPR-R-S. This improvement stems from the knowledge learned via structure-aware distillation during multimodal training. The benefit is particularly evident on several single-chip radar sequences (BS, SON, SQ).
IV-F Study on Viewpoint Rotation Invariance